What is the capacitance of the empty shelves if they have area of 1.00x10 m and are 0.200 m apart?


1
Expert's answer
2022-03-16T09:42:05-0400

The capacitance of the parallel plates capacitor

"C=\frac{\epsilon_0 A}{d}""C=\frac{8.85*10^{-12}*10.0}{0.200}=4.43*10^{-10}\:\rm F"